A truck starts at Point A and drives South three miles. Then it turns left at Point B and drives East four miles to Point C. How many miles is the truck at Point C from Point A if it were to drive directly from point A to Point C?
Answer by stanbon(48558)

A truck starts at Point A and drives South three miles. Then it turns left at Point B and drives East four miles to Point C. How many miles is the truck at Point C from Point A if it were to drive directly from point A to Point C?
------------
distance = sqrt[3^2 + 4^2] = sqrt[9+16] = sqrt(25) = 5 miles
====================================================================
Cheers,
Stan H